# Benford behavior resulting from stick and box fragmentation processes

### v2

@@ 3.1. Preliminary
To find the PDF of $Y^{(N)}$, we need the joint PDF of $Z_{(m+1-i)}^{(N)}$ for all $1\leq i\leq d$. This is a standard question in order statistics.
Proposition 3.3.
[DL] For i.i.d. random variables $X_{1},\dots,X_{k}$ with PDF $f(x)$ and CDF $F(x)$, the joint PDF of $X_{(i)},\dots,X_{(k)}$ where $1\leq i\leq k$ is given by
$f_{X_{(i)},\dots,X_{(k)}}(x_{i},\dots,x_{k})\ =\ C_{k}^{i}f(x_{i})\cdots f(x_{k})(F(x_{i}))^{i-1},$ (3.6)
where $C_{k}^{i}:=k!/(i-1)!$.
Thus, in our case, the joint PDF of $Z_{(m+1-i)}^{(N)}$ for all $1\leq i\leq d$ is
$f_{Z_{(m+1-d)},\dots,Z_{(m)}}(z_{m+1-d},\dots,z_{m})$
$\ =\ C_{m}^{m+1-d}\left(\Phi(z_{m+1-d})+B(z_{m+1-d})\right)^{m-d}\prod_{i=1}^{d}\left(\varphi(z_{m+i-d})+A(z_{m+i-d})\right).$ (3.7)
By binomial expansion,
$\left(\Phi(z_{m+1-d})+B(z_{m+1-d})\right)^{m-d}\ =\ \sum_{j=0}^{m-d}\binom{m-d}{j}(\Phi(z_{m+1-d}))^{m-d-j}(B(z_{m+1-d}))^{j}.$ (3.8)
Note that $B(z_{m+1-d})=O(N^{-\delta})$ and $0\leq\Phi(x)\leq 1$, which we shall constantly use without reference for the remainder of the paper. Then any term where $j>1$ is at most of the same magnitude as $B(z_{m+1-d})$. We have
$\left(\Phi(z_{m+1-d})+B(z_{m+1-d})\right)^{m-d}\ =\ \left(\Phi(z_{m+1-d})\right)^{m-d}+O(B(z_{m+1-d})).$ (3.9)
@@
We make a slight abuse of notation here to write the error term above as $B(z_{m+1-d})$, which does not change the order of magnitude of the error term. Hence,
$f_{Z_{(m+1-d)},\dots,Z_{(m)}}(z_{m+1-d},\dots,z_{m})$
$\ =\ C_{m}^{m+1-d}\left(\Phi(z_{m+1-d})^{m-d}+B(z_{m+1-d})\right)\prod_{i=1}^{d}\left(\varphi(z_{m+i-d})+A(z_{m+i-d})\right).$ (3.10)
3.2. PDF of $Y^{(N)}$
@@
In this section, we find the PDF of $Y^{(N)}$ as defined in (3.1). The CDF $F_{Y^{(N)}}(y)$ of $Y^{(N)}$ is given by integrating over the appropriate region in $\mathbb{R}^{d}$, that is, over all the values $(z_{m+1-d},\dots,z_{m})$ of $(Z_{(m+1-d)}^{(N)},\dots,Z_{(m)}^{(N)})$ that sums to $y$. We know that $Z_{(m+1-d)}^{(N)},\dots Z_{(m)}^{(N)}$ are not independent, i.e., they have to satisfy $Z_{(m+1-d)}^{(N)}\leq\cdots\leq Z_{(m)}^{(N)}$. In general, we know that any $1\leq j\leq d$, $z_{m+j-d}$ can take value from $z_{m+(j-1)-d}$ to $(y-\sum_{i=1}^{j-1}z_{m+i-d})/(d-j+1)$. The only exception is $z_{m+1-d}$, whose upper bound is already correctly stated but the lower bound is $C\sqrt{N}$. Hence, we have
@@
$\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z_{(m+1-d)},\dots,Z_{(m)}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d},$ (3.11)
@@
$\prod_{j=k}^{d}\int_{z_{j}}^{y_{j}}\ :=\ \int_{z_{k}}^{y_{k}}\cdots\int_{z_{j}}^{y_{j}}\cdots\int_{z_{d}}^{y_{d}}.$ (3.12)
@@
$m_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})\ :=\ C_{m}^{m+1-d}\Phi(z_{m+1-d})^{m-d}\prod_{i=1}^{d}\varphi(z_{m+i-d}).$ (3.13)
We also denote the error term of the joint PDF (3.1) by $e_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})$. Let $\mathcal{M}_{Y^{(N)}}(y)$ and $\mathcal{E}_{Y^{(N)}}(y)$ be the main term and the error term of $F_{Y^{(N)}}(y)$. Then
$F_{Y^{(N)}}(y)$ $\ =\ \mathcal{M}_{Y^{(N)}}(y)+\mathcal{E}_{Y^{(N)}}(y)$
$\mathcal{M}_{Y^{(N)}}(y)$ $\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}m_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}$
$\mathcal{E}_{Y^{(N)}}(y)$ $\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}e_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}.$ (3.14)
@@
$f_{Y^{(N)}}(y)\ =\ \frac{d}{dy}F_{Y^{(N)}}(y)\ =\ \frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)+\frac{d}{dy}\mathcal{E}_{Y^{(N)}}(y).$ (3.15)
@@
To differentiate $\mathcal{M}_{Y^{(N)}}(y)$ and $\mathcal{E}_{Y^{(N)}}(y)$, we apply the Leibniz integral rule for differentiate under the integral sign.
Lemma 3.4.
@@
[F] Suppose that $f(x,t)$ is a function such that both $f(x,t)$ and $df(x,t)/dx$ are continuous in $t$ and $x$ in the $xt$-plane, and $a(x)$ and $b(x)$ are also continuously differentiable. Then
$\frac{d}{dx}\int_{a(x)}^{b(x)}f(x,t)dt\ =\ f(x,b(x))\cdot\frac{d}{dx}b(x)-f(x,a(x))\cdot\frac{d}{dx}a(x)+\int_{a(x)}^{b(x)}\frac{\partial}{\partial x}f(x,t)dt.$ (3.16)
@@
For the rest of this paper, one can easily check the regularity conditions on $f(x,t)$, $a(x)$, and $b(x)$ each time we apply the Leibniz integral rule, so there will not be explicit mention of the regularity conditions again. We first make some observations on the effect of $d/dy$ on
$\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}g(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}$ (3.17)
@@
for any Riemann integrable function $g(z_{m+1-d},\dots,z_{m})$. One can check that each application of the Leibniz integral rule to (3.17) gives two integrals, one of which has the lower bound of the integration coinciding with the upper bound of the integration and thus equals to 0. As a result, we are always left with one integral that interchanges the order of the differentiation and the integration signs. Eventually, (3.17) becomes
@@
$\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\frac{\partial}{\partial y}\left(\int_{z_{m-1}}^{y-\sum_{i=1}^{d-1}z_{m+i-d}}g(z_{m+1-d},\dots,z_{m})dz_{m}\right)dz_{m-1}\cdots dz_{m+1-d}$
@@
$\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}g\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\cdots dz_{m+1-d}.$ (3.18)
Applying (3.2) to $d\mathcal{M}_{Y^{(N)}}(y)/dy$ and $d\mathcal{E}_{Y^{(N)}}(y)/dy$, where $\mathcal{M}_{Y^{(N)}}(y)$ and $\mathcal{E}_{Y^{(N)}}(y)$ are as defined in (3.2), we have
@@
$\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)$ $\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}m_{Y^{(N)}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}$
$\quad\quad\cdots dz_{m+1-d}$

### v3

@@ 3.1 Preliminary
To find the PDF of $Y^{(N)}$, we need the joint PDF of $Z_{(m+1-i)}^{(N)}$ for all $1\leq i\leq d$, which is a standard question in order statistics.
Proposition 5.
[9] For i.i.d. random variables $X_{1},\dots,X_{n}$ with PDF $f(x)$ and CDF $F(x)$, the joint PDF of $X_{(i)},\dots,X_{(n)}$ is given by
$f_{X_{(i)},\dots,X_{(n)}}(x_{i},\dots,x_{n})\ =\ C_{n}^{i}f(x_{i})\cdots f(x_{n})(F(x_{i}))^{i-1},$ (53)
where $1\leq i\leq n$ and $C_{n}^{i}:=n!/(i-1)!$.
Since $Z_{i}^{(N)}$’s are i.i.d., then to find the joint PDF of $Z_{(m+-i)}^{(N)}$, it suffices to find the PDF and CDF of $Z_{i}^{(N)}$’s, which converge to that of the standard Gaussian distribution by CLT. However, to gain a sharp estimate on the distribution of $Y^{(N)}$, we would need to quantify the error terms of the PDF and the CDF, as follows.
To begin, we prove the following Riemann-Lebesgue lemma type bound on the characteristic function of a random variable.
Lemma 3.
Suppose that $f(x)$ is the PDF of a random variable $X$, where $f(x)$ is differentiable with $\sup_{x\in\mathbb{R}}|f^{\prime}(x)|<\infty$ and $X$ has finite mean and nonzero, finite variance. Then the characteristic function
$\psi(t)\ :=\ \int_{-\infty}^{\infty}f(x)e^{itx}dx$ (54)
satisfies $\psi(t)=O(t^{-2/5})$ for any $\epsilon>0$.
Proof.
We know that $e^{itx}=\cos(tx)+i\sin(tx)$. Hence, it suffices to show that
$\int_{-\infty}^{\infty}f(x)\cos(tx)dx\ =\ O(t^{-2/5})\quad\textup{and}\quad\int_{-\infty}^{\infty}f(x)\sin(tx)dx\ =\ O(t^{-2/5}).$ (55)
The proof for the two bounds are similar, and we only show it for $\int_{-\infty}^{\infty}f(x)\cos(tx)dx$. We first estimate the tail of the integral.
$\left|\int_{|x|\geq t^{1/5}}f(x)\cos(tx)dx\right|$ $\ \leq\ \int_{|x|\geq t^{1/5}}f(x)dx$
$\ =\ \mathbb{P}\left(|X|\geq t^{1/5}\right)$
$\ \leq\ \mathbb{P}\left(|X-\mathbb{E}[X]|\geq t^{1/5}-\mathbb{E}[X]\right)+\mathbb{P}\left(|X-\mathbb{E}[X]|\geq t^{1/5}+\mathbb{E}[X]\right)$
$\ \leq\ \frac{\textup{Var}[X]}{(t^{1/5}-\mathbb{E}[X])^{2}}+\frac{\textup{Var}[X]}{(t^{1/5}+\mathbb{E}[X])^{2}}\ =\ O(t^{-2/5}),$ (56)
where the last line follows from Chebyshev’s inequality, which requires $X$ to have finite mean and nonzero, finite variance. Hence, we can restrict the domain of integration to $t^{1/5},t^{1/5}]$. We know that if $g(x)$ is differentiable on $[a,b]$ with $\sup_{x\in[a,b]}|g^{\prime}(x)|=M<\infty$ [14], then for any Riemann sum $S=\sum_{k=1}^{n}g(x_{k}^{*})(x_{k}-x_{k-1})$,
$\left|\int_{-\infty}^{\infty}g(x)dx-S\right|\ \leq\ \frac{M(b-a)\delta_{\textup{max}}}{2},$ (57)
where $\delta_{max}=\max_{k}(x_{k}-x_{k-1})$. Now consider a Riemann sum $S_{t}=\sum_{k=1}^{n}f(x_{k}^{*})(x_{k}-x_{k-1})$ with the partition $x_{k}-x_{k-1}=t^{-3/5}$ for all $k$. Define the function $f_{t}$ on $t^{1/5},t^{1/5}]$ such that $f_{t}(x):=f(x_{k}^{*})$ for all $x\in[x_{k-1},x_{k})$ and for every $k$. Since $f$ is nonnegative, then $f_{t}$ is also nonnegative. Hence,
$\left|\int_{-t^{1/5}}^{t^{1/5}}f(x)\cos(tx)dx-\int_{-t^{1/5}}^{t^{1/5}}f_{t}(x)\cos(tx)dx\right|$ $\ \leq\ \left|\int_{-t^{1/5}}^{t^{1/5}}f(x)dx-S_{t}\right|$
$\ \leq\ \frac{M_{t}\cdot(2t^{1/5})\cdot t^{-3/5}}{2}\ =\ O(t^{-2/5}),$ (58)
where $M_{t}:=\sup_{x\in[-t^{1/5},t^{1/5}]}|f^{\prime}(x)|<\infty$. Moreover, we have
$\int_{-t^{1/5}}^{t^{1/5}}f_{t}(x)\cos(tx)dx$ $\ =\ \int_{-t^{1/5}}^{t^{1/5}}\sum_{k=1}^{n}f(x_{k}^{*})\chi_{[x_{k-1},x_{k})}(x)\cos(tx)dx$
$\ =\ \sum_{k=1}^{n}f(x_{k}^{*})\int_{x_{k-1}}^{x_{k}}\cos(tx)dx$
$\ =\ \frac{1}{t}\sum_{k=1}^{n}f(x_{k}^{*})(\sin(tx_{k})-\sin(tx_{k-1}))$
$\ =\ O\left(\frac{1}{t}\sum_{k=1}^{n}f(x_{k}^{*})\right).$ (59)
Since $S_{t}=t^{-3/5}\sum_{k=1}^{n}f(x_{k}^{*})$ and that the Riemann sum is finite for large $t$, then
$\int_{-t^{1/5}}^{t^{1/5}}f_{t}(x)\cos(tx)dx$ $\ =\ O(t^{-2/5}).$ (60)
Thus, combining (3.1), (3.1), and (60), we have
$\left|\int_{-\infty}^{\infty}f(x)\cos(tx)dx\right|$
$\ \leq\ \left|\int_{|x|\geq t^{1/5}}f(x)\cos(tx)dx\right|+\left|\int_{-t^{1/5}}^{t^{1}/5}f(x)\cos(tx)dx-\int_{-t^{1/5}}^{t^{1/5}}f_{t}(x)\cos(tx)dx\right|$
$\quad\quad+\left|\int_{-t^{1/5}}^{t^{1/5}}f_{t}(x)\cos(tx)dx\right|\ =\ O(t^{-2/5}).$ (61)
Lemma 4.
Let $Z_{i}^{(N)}\sim Z^{(N)}$. For any $0<\epsilon<1/2$, the PDF $f_{Z^{(N)}}$ and the CDF $F_{Z^{(N)}}$ of $Z^{(N)}$ satisfy
$f_{Z^{(N)}}(z)\ =\ \varphi(z)+O(N^{-3/2+3\epsilon})$ (62)
$F_{Z^{(N)}}(z)\ =\ \Phi(z)+O(N^{-1/2}),$ (63)
where $\varphi$ and $\Phi$ are the PDF and the CDF of the standard Gaussian distribution.
Proof.
We first prove (62). Let the characteristic function of $W:=(\log_{B}(P)-\mu_{P})/\sigma_{P}$ be $\psi_{W}(t)$, which has mean 0, variance $1$, and finite moments by Condition 2. Hence,
$\psi_{W}(t)$ $\ =\ \sum_{k=0}^{\infty}\frac{\mathbb{E}[W^{k}](it)^{k}}{k!}$
= $\ =\ 1-\frac{t^{2}}{2}+O(t^{3})$ (64)
for sufficiently small $t$. Let the characteristic function of $Z^{(N)}$ be $\psi_{Z^{(N)}}(t)$. By properties of characteristic functions,
$\psi_{Z^{(N)}}(t)$ $\ =\ (\psi_{W}(t/\sqrt{N}))^{N}$
$\ =\ \left(1-\frac{t^{2}}{2N}+O\left(\frac{t^{3}}{N^{3/2}}\right)\right)^{N}$
$\ =\ \textup{exp}\left[N\log\left(1-\frac{t^{2}}{2N}+O\left(\frac{t^{3}}{N^{3/2}}\right)\right)\right]$
$\ =\ \textup{exp}\left[N\left(\log\left(1-\frac{t^{2}}{2N}\right)+O\left(\frac{1}{1-t^{3}/N^{3/2}}\cdot\frac{t^{3}}{N^{3/2}}\right)\right)\right]$
$\ =\ \textup{exp}\left[N\left(-\frac{t^{2}}{2N}+O\left(\frac{t^{4}}{N^{2}}\right)+O\left(\frac{t^{3}}{N^{3/2}}\right)\right)\right]$
$\ =\ e^{-t^{2}/2}\left(1+O\left(\frac{t^{3}}{N^{3/2}}\right)\right),$ (65)
where in the last two lines, we make use of the Taylor series expansions of $\log(1+x)$ and $e^{x}$. Note that the above estimate holds for $|t|\leq N^{\epsilon}$ for any $0<\epsilon<1/2$. By Fourier inversion, the PDF of $Z^{(N)}$ is
$\frac{1}{2\pi}\int_{-\infty}^{\infty}\psi_{Z^{(N)}}(t)e^{-itx}dt.$ (66)
We shall split the integral above into several parts and provide an estimate for each. We first estimate the integral for $|t|\leq N^{\epsilon}$. By (3.1),
$\frac{1}{2\pi}\int_{|t|\leq N^{\epsilon}}\psi_{Z^{(N)}}(t)e^{-itx}dt\ =\ \varphi(x)+O(N^{-3/2+3\epsilon}).$ (67)
By Lemma 3, we know that $\psi_{W}(t)=O(t^{-2/5})$. Let $E$ be a constant such that $E\geq(\pi/4)^{2/5}$ and $|\psi_{W}(t)|\leq Et^{-2/5}$ for sufficiently large $t$. Then, for any $\epsilon_{0}>0$, we have
$\left|\frac{1}{2\pi}\int_{|t|\geq(E+\epsilon_{0})^{5/2}\sqrt{N}}\psi_{Z^{(N)}}(t)e^{-itx}dt\right|$ $\ \ll\ \int_{t\geq(E+\epsilon_{0})^{5/2}\sqrt{N}}\left(\frac{E}{(t/\sqrt{N})^{2/5}}\right)^{N}dt$
$\quad\ \ll\ \sqrt{N}E^{N}\int_{t\geq(E+\epsilon_{0})^{5/2}}t^{-2N/5}dt$
$\quad\ \ll\ \frac{\sqrt{N}}{N}\cdot\frac{E^{N}}{(E+\epsilon_{0})^{N}},$ (68)
which is of exponential decay in $N$. By [7] and Condition 3, for $|t|\geq\pi/4$, we have
$|\psi_{W}(t)|\ <\ 1-\frac{c_{1}}{m^{2}},$ (69)
where $m$ is the median of $W$’s which is finite by Condition 3, and $c_{1}>0$ is some constant. Hence,
$\left|\frac{1}{2\pi}\int_{\pi\sqrt{N}/4\leq|t|\leq(E+\epsilon_{0})^{5/2}\sqrt{N}}\psi_{Z^{(N)}}(t)e^{-itx}dt\right|\ \ll\ \sqrt{N}\left(1-\frac{c_{1}}{m^{2}}\right)^{N},$ (70)
which is also of exponential decay in $N$. By [7], when $0<t<\pi/4$, we have
$|\psi_{W}(t)|\ <\ 1-\frac{c_{2}t^{2}}{m},$ (71)
where $m$ is again the median of $W_{i}$’s and $c_{2}>0$ is some constant. Hence,
$\left|\frac{1}{2\pi}\int_{N^{\epsilon}\leq|t|\leq\pi\sqrt{N}/4}\psi_{Z^{(N)}}(t)e^{-itx}dt\right|$ $\ \ll\ \sqrt{N}\left(1-\frac{c_{2}}{m^{2}}N^{2\epsilon-1}\right)^{N}$
$\ \ll\ \sqrt{N}e^{-2c_{2}N^{2\epsilon}/m^{2}},$ (72)
which also decays exponential in $N$. Combining (67), (3.1), (70), (3.1), we have
$f_{Z^{(N)}}(x)\ =\ \frac{1}{2\pi}\int_{-\infty}^{\infty}\psi_{Z^{(N)}}(t)e^{-itx}dt\ =\ \varphi(x)+O(N^{-3/2+3\epsilon}),$ (73)
which is the desired PDF for $Z^{(N)}$. Moreover, by Berry-Esseen theorem [4, 11], we have that for all $N$ and $x$,
$F_{Z^{(N)}}(x)\ =\ \Phi(x)+O(N^{-1/2}),$ (74)
which is the desired estimate for the CDF of $Z^{(N)}$. ∎
As we shall see for the rest of this section, it suffices for the error term of $f_{Z^{(N)}}$ to be $O(N^{-1/2-\delta})$ and the error term of $F_{Z^{(N)}}$ to be $O(N^{-\delta})$ for some $\delta>0$. Hence, the estimates of Lemma 4 are more than enough for this purpose. Now, we go back to (53) given in Proposition 5. In our case, the joint PDF of $Z_{(m+1-i)}^{(N)}$ for all $1\leq i\leq d$ is
$f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)}}(z_{m+1-d},\dots,z_{m})$
$\quad\ =\ C_{m}^{m+1-d}\left(\Phi(z_{m+1-d})+B(z_{m+1-d})\right)^{m-d}\prod_{i=1}^{d}\left(\varphi(z_{m+i-d})+A(z_{m+i-d})\right),$ (75)
where $A(x)=O(N^{-1/2-\delta})$ and $B(x)=O(N^{-\delta})$. By binomial expansion,
$\left(\Phi(z_{m+1-d})+B(z_{m+1-d})\right)^{m-d}\ =\ \sum_{j=0}^{m-d}\binom{m-d}{j}(\Phi(z_{m+1-d}))^{m-d-j}(B(z_{m+1-d}))^{j}.$ (76)
Since $\Phi(x)$ is a CDF, then $0\leq\Phi(x)\leq 1$, a fact which we shall use without explicit mention for the remainder of the paper. Then, any term in (76) with $j>1$ is at most of the same order as $B(z_{m+1-d})$. Hence,
$\left(\Phi(z_{m+1-d})+B(z_{m+1-d})\right)^{m-d}\ =\ \left(\Phi(z_{m+1-d})\right)^{m-d}+O(B(z_{m+1-d})).$ (77)
@@
We make a slight abuse of notation here to write the error term above as $B(z_{m+1-d})$, which does not change the order of the error term. So then we have
$f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)}}(z_{m+1-d},\dots,z_{m})$
$\quad\ =\ C_{m}^{m+1-d}\left(\Phi(z_{m+1-d})^{m-d}+B(z_{m+1-d})\right)\prod_{i=1}^{d}\left(\varphi(z_{m+i-d})+A(z_{m+i-d})\right).$ (78)
3.2 PDF of $Y^{(N)}$
@@
In this section, we find the PDF of $Y^{(N)}$. The CDF $F_{Y^{(N)}}(y)$ of $Y^{(N)}$ is given by integrating over the appropriate region in $\mathbb{R}^{d}$, that is, over all the values $(z_{m+1-d},\dots,z_{m})$ of $(Z_{(m+1-d)}^{(N)},\dots,Z_{(m)}^{(N)})$ that sums to $y$. We know that $Z_{(m+1-d)}^{(N)},\dots Z_{(m)}^{(N)}$ are not independent, i.e., they have to satisfy $Z_{(m+1-d)}^{(N)}\leq\cdots\leq Z_{(m)}^{(N)}$. In general, we know that any $1\leq j\leq d$, $z_{m+j-d}$ can take value from $z_{m+(j-1)-d}$ to $(y-\sum_{i=1}^{j-1}z_{m+i-d})/(d-j+1)$. The only exception is $z_{m+1-d}$, whose upper bound is already correctly stated but the lower bound is -$\infty$. Hence, we have
@@
$\quad\ =\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d},$ (79)
@@
$\prod_{j=k}^{d}\int_{z_{j}}^{y_{j}}\ :=\ \int_{z_{k}}^{y_{k}}\cdots\int_{z_{j}}^{y_{j}}\cdots\int_{z_{d}}^{y_{d}}.$ (80)
@@
$f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ main}}(z_{m+1-d},\dots,z_{m})\ :=\ C_{m}^{m+1-d}\Phi(z_{m+1-d})^{m-d}\prod_{i=1}^{d}\varphi(z_{m+i-d}).$ (81)
We also denote the error term of the joint PDF (3.1) by $f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ error}}(z_{m+1-d},\dots,z_{m})$. Let $F_{Y^{(N)},\textup{ main}}(y)$ and $F_{Y^{(N)},\textup{ error}}(y)$ be the main term and the error term of $F_{Y^{(N)}}(y)$. Then
$F_{Y^{(N)}}(y)$ $\ =\ F_{Y^{(N)},\textup{ main}}(y)+F_{Y^{(N)},\textup{ error}}(y)$
$F_{Y^{(N)},\textup{ main}}(y)$ $\ =\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ main}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}$
$F_{Y^{(N)},\textup{ error}}(y)$ $\ =\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ error}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}.$ (82)
@@
$f_{Y^{(N)}}(y)\ =\ \frac{d}{dy}F_{Y^{(N)}}(y)\ =\ \frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)+\frac{d}{dy}F_{Y^{(N)},\textup{ error}}(y).$ (83)
@@
To differentiate $F_{Y^{(N)}(y),\textup{ main}}$ and $F_{Y^{(N)},\textup{ error}}(y)$, we apply the Leibniz integral rule for differentiate under the integral sign.
Lemma 5.
@@
[12] Suppose that $f(x,t)$ is a function such that both $f(x,t)$ and $df(x,t)/dx$ are continuous in $t$ and $x$ in the $xt$-plane, and $a(x)$ and $b(x)$ are also continuously differentiable. Then
$\frac{d}{dx}\int_{a(x)}^{b(x)}f(x,t)dt\ =\ f(x,b(x))\cdot\frac{d}{dx}b(x)-f(x,a(x))\cdot\frac{d}{dx}a(x)+\int_{a(x)}^{b(x)}\frac{\partial}{\partial x}f(x,t)dt.$ (84)
@@
For the rest of this paper, one can easily check the regularity conditions on $f(x,t)$, $a(x)$, and $b(x)$ each time we apply the Leibniz integral rule, so there will not be any explicit mention of the regularity conditions again. We first make some observations on the effect of $d/dy$ on
$\int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}g(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}$ (85)
@@
for any continuous function $g(z_{m+1-d},\dots,z_{m})$. One can check that each application of the Leibniz integral rule to (85) would only yield one integral, because for the other integral, the lower bound of the integration always coincides with the upper bound of the integration and thus equals to 0. As a result, each application of $d/dy$ here merely interchanges the order of the differentiation and the integral signs. Eventually, (85) becomes
@@
$\int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\frac{\partial}{\partial y}\left(\int_{z_{m-1}}^{y-\sum_{i=1}^{d-1}z_{m+i-d}}g(z_{m+1-d},\dots,z_{m})dz_{m}\right)dz_{m-1}\cdots dz_{m+1-d}$
@@
$\quad\ =\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}g\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\cdots dz_{m+1-d}.$ (86)
Applying (3.2) to $dF_{Y^{(N)}}/dy$, we have that the PDF of $Y^{(N)}$ is
$\frac{d}{dy}F_{Y^{(N)}}(y)\ =\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
$\quad\quad\quad\quad\quad\quad\quad dz_{m-1}\cdots dz_{m+1-d}.$ (87)
Applying (3.2) to $dF_{Y^{(N)},\textup{ main}}(y)/dy$ and $dF_{Y^{(N)},\textup{ error}}(y)/dy$, we have
$\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)$
@@
$\quad\ =\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ main}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
$\quad\quad\quad dz_{m-1}\cdots dz_{m+1-d}$
$\frac{d}{dy}F_{Y^{(N)},\textup{ error}}(y)$